OpenStudy (anonymous):

how do you find the solutions to \[\cos 2\theta-\sin \theta=1\]?

OpenStudy (jamesj):

Remember that cos 2x = cos^2 x - sin^2 x = 1 - 2 sin^2 x Hence your equation can be written as a quadratic equation in sin

OpenStudy (anonymous):

ok so i get 2sin^2theta+sintheta-2. how do i solve to find solutions from here?

OpenStudy (jamesj):

Write, for instance, x = sin theta. Then you have a quadratic in x. Solve that quadratic in x. And then find theta such that sin theta = x.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

hello again, do you know how to prove the equation, \[(\sin 3\theta \div \sin \theta)-(\cos3\theta \div \cos \theta)=2\], is an idenity?